OpenStudy (anonymous):

The circuit elements have the following values: Vin=27V, R1=5Ω, R2=7Ω, R3=12Ω, A=6 and B=0.1

OpenStudy (anonymous):

i am not sure about the above values..i still need to check :p

OpenStudy (anonymous):

Rth????

OpenStudy (anonymous):

rth=vth/isc

OpenStudy (anonymous):

isc formula plzzz gyuz

OpenStudy (anonymous):

@1431mahi i1=Vo/(R3-(B*R1*R3) Vo=(Vin*(R3-B*R1*R3))/(R3-(B*R1*R2)+R1+R2+(a*R2)) Vth=Vo Rth=Vth/Isc put Isc=0.25

OpenStudy (anonymous):

Isc = 0.25 . . how??

OpenStudy (anonymous):

Isc is the short circuit current..... short the independent voltage source and find the current entering from the output port...

OpenStudy (anonymous):

@neelima and what about dependent sources

OpenStudy (anonymous):

Rth=10.8 vo=2.7

OpenStudy (anonymous):

@akash97 how you have Rth . . what are the circuit parameters

OpenStudy (anonymous):

@ali_amjad : deondnt sources cant be replaced by der intrnal resistances

OpenStudy (anonymous):

the following values: Vin=15V, R1=0.5Ω, R2=1.5Ω, R3=5Ω, A=2 and B=1.5. my answer : 3 3 2.4 Rth ?? Isc??

OpenStudy (anonymous):

@akash97 any hint now??

OpenStudy (anonymous):

@ali_amjad : short ckt d o/p terminal n get Isc

OpenStudy (anonymous):

Shorted it and also R3 will have no effect now??

OpenStudy (anonymous):

@ali_amjad : find currnt through dt mesh

OpenStudy (anonymous):

there two masehs?? one containg R3 and Bvb and second containg Ava and R1 R2?? correct